A 140-lb person is standing on a bathroom scale in an elevator as the elevator accelerates downward 1.2 m/s2 What does the bathroom scale read?

Answer is 120 lbs.

Please show step by step solution to this problem.
Answer #1

1 lb = 0.4536 kg

140 lb = 63.504 kg

now, as the elevator moving downwards

W = m ( g - a)

W = 63.504* ( 9.8 + 1.2)

W = 546.134 N

W = 546.134 / ( 9.8 *0.467)

W = 119.332 lb ≈ 120 lb

=======

Comment before rate in case any doubt, will reply for sure.. goodluck